# A new approach to principal–agent problems with volatility control

### v1

@@ 3. $(iii)$
Usually, 1 cannot be solved directly: the standard approach is to determine an appropriate form of contracts which simplifies both the agent and the principal’s problems, and such that the restriction to this form of contracts is without loss of generality.
Note that the previous setting and assumptions are relatively standard in continuous-time principal–agent problems, and in particular coincide with the framework described in [9]. Indeed, as the goal of this paper is to provide a new approach to solve such principal–agent problems, which in particular requires to show that this approach allows to achieve the same results obtained by Cvitanić, Possamaï, and Touzi (2018) [9], the decision was made to closely follow their original framework. It is nevertheless possible that some assumptions could be relaxed, or at least modified to consider alternative frameworks. For example, in the case of a principal–agent problem with non-continuous output processes (and many agents), one can refer to the recent paper by Hernández Santibáñez (2022) [13] for more general assumptions. It should be noted, however, that this paper focuses on control processes that can be represented through equivalent changes of measures, thus a priori excluding volatility control.
@@ 2.3 The existing approach, via 2BSDEs
In this section, we describe briefly and rather informally the general approach developed by Cvitanić, Possamaï, and Touzi (2018) [9], as this will help in comparing it with our new approach, described later in Section 3.
@@
The general approach developed in [9] involves first identifying a class of contracts, offered by the principal, that are revealing, in the sense that the agent’s optimal response to the contract can easily be computed. More precisely, in the framework under consideration here, we first define the appropriate agent’s hamiltonian, for any $(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d}$,
@@
We then introduce the relevant form of contracts, namely $\xi=Y_{T}^{y_{0},Z,\Gamma}$, where the process $Y^{y_{0},Z,\Gamma}$ is defined as the solution to an ordinary differential equation (ODE for short) with random coefficients,
@@
for some processes $(Z,\Gamma)$ satisfying appropriate integrability conditions. More precisely, for any $\mathbb{F}$-predictable $\mathbb{R}^{d}$-valued process $Z$ and any $\mathbb{F}$-optional $\mathbb{R}$-valued process $Y$, we are led to consider the following norms,
@@ 2. $(ii)$
Then, the relevant contracts $\xi\in\Xi$ are those that admit the representation $\xi=Y_{T}^{y_{0},Z,\Gamma}$, where the pair $(Z,\Gamma)\in{\cal V}$ as well as the constant $y_{0}\in\mathbb{R}$ in (2.11) will be optimally chosen by the principal, in order to maximise her expected utility, but also to ensure the participation of the agent.
@@
Contracts of the form (2.11) are called revealing, as it is straightforward to compute the agent’s best response to such contracts. Indeed, relying on Ito’s formula, it is proved in [9, Proposition 3.3] that, given a contract of the form $\xi=Y_{T}^{y_{0},Z,\Gamma}$, $(y_{0},Z,\Gamma)\in\mathbb{R}\times{\cal V}$, any optimal response $\mathbb{P}^{\star}\in{\cal P}^{\star}(\xi)$ is such that
@@
and remark that the agent’s best response at time $t\in[0,T]$ to a contract $\xi=Y_{T}^{y_{0},Z,\Gamma}$ is therefore given by $\nu^{\mathbb{P}^{\star}}_{t}:=u^{\star}(t,X,Y_{t}^{y_{0},Z,\Gamma},Z_{t},\Gamma_{t}),\;\mathbb{P}^{\star}$–a.s. One can also prove that the value of the agent’s optimal control problem associated to such contract $\xi$ of the form (2.11) boils down to $y_{0}$, namely $V_{\rm A}(Y_{T}^{y_{0},Z,\Gamma})=y_{0}$. Consequently, to ensure the agent’s participation in the contractual agreement, it is clear that the principal must choose $y_{0}\geq R_{\rm A}$.
@@
$\mathrm{d}Y_{t}^{y_{0},Z,\Gamma}=$ $\ \Big{(}c\big{(}t,X,\nu_{t}^{\mathbb{P}^{\star}}\big{)}+Y_{t}^{y_{0},Z,\Gamma}k_{\rm A}\big{(}t,X,\nu_{t}^{\mathbb{P}^{\star}}\big{)}\Big{)}\mathrm{d}t+Z_{t}\cdot\sigma\big{(}t,X,\nu_{t}^{\mathbb{P}^{\star}}\big{)}\mathrm{d}W_{t},\quad t\in[0,T].$ (2.16)
@@
$\mathrm{d}Y_{t}$ $=\Big{(}c\big{(}t,X,\nu^{\mathbb{P}^{\star}}_{t}\big{)}+Y_{t}k_{\rm A}\big{(}t,X,\nu^{\mathbb{P}^{\star}}_{t}\big{)}\Big{)}\mathrm{d}t+Z_{t}\cdot\sigma\big{(}t,X,\nu^{\mathbb{P}^{\star}}_{t}\big{)}\mathrm{d}W_{t},\quad t\in[0,T],$ (2.17b)
coupled through the agent’s optimal effort $\nu^{\mathbb{P}^{\star}}_{t}:=u^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t})$ under $\mathbb{P}^{\star}\in{\cal P}^{\star}(Y_{T})$, where the function $u^{\star}$ is defined as a maximiser of the Hamiltonian through (2.13).
In a Markovian setting, one can solve this stochastic control problem as usual, namely by writing the principal’s Hamiltonian as well as the Hamilton-Jacobi-Bellman (HJB for short) equation satisfied by the value function $\underline{V}_{\rm P}(y_{0})$ (see [9, Theorem 3.9]), and finally optimise it over $y_{0}\geq R_{\rm A}$.
More importantly for our purpose in this paper, one can already remark that the value of 2 should be lower than the value in 1, since considering contracts $\xi$ of the form (2.11) is a restriction of the original set $\Xi$ of admissible contracts. This result can be found in [9, Proposition 3.4], and is a direct consequence of [9, Proposition 3.3] mentioned above. For further reference, this result is stated in the lemma below.
@@
The original value defined by 1 is greater than the value in 2, namely $V_{\rm P}\geq\widetilde{V}_{\rm P}$.
@@
To summarise, when restricting to contracts $\xi\in\Xi$ of the form (2.11), one can first easily determine the agent’s optimal response to such contract, namely by computing the maximiser of his Hamiltonian. Then, still restricting to such revealing contract, the principal’s problem boils down to a more standard stochastic problem, namely 2. These two steps are therefore quite straightforward using standard tools from stochastic control theory. However, the last and crucial step requires to prove that the restriction to such revealing contracts is without loss of generality. More precisely, one need to argue that the value of the original principal’s problem, namely $V_{\rm P}$ recalled in 1, actually coincides with the value when restricting to revealing contracts, i.e. $\widetilde{V}_{\rm P}$ in 2. While it is clear that $V_{\rm P}\geq\widetilde{V}_{\rm P}$, as mentioned in Lemma 2.5 above, the main mathematical challenge is to show the converse inequality, namely $V_{\rm P}\leq\widetilde{V}_{\rm P}$. This result, stated in [9, Theorem 3.6], relies on the sophisticated theory of second-order BSDEs. We deliberately omit to develop further this step here, as it is precisely the part we can avoid with the new approach. It nevertheless seems essential to stress that this is the most difficult stage, which probably prevents a wider application of principal–agent problems with volatility control, but also hampers natural extensions of these problems, as discussed in introduction.
@@ 3 A new approach
In this section, we first start by formalising in Section 3.1 a ‘first-best’ alternative problem, in which the principal, in addition to offering a contract, also controls the quadratic variation of the output process, through its density with respect to the Lebesgue measure, denoted by $\Sigma$. We then provide in Section 3.2 a straightforward resolution approach for this alternative problem, relying only on the theory of BSDEs. We will finally prove in Section 3.3 that this reformulation actually leads to the exact same solution compared to the original problem considered in Section 2. In other words, from the principal’s point of view, optimising directly the quadratic variation through its density process $\Sigma$ is strictly equivalent to choosing $\Gamma$ in the contract form (2.11).
@@ 3.1 The ‘first-best’ formulation
As already mentioned, the approach developed by Cvitanić, Possamaï, and Touzi (2018) [9] mostly relies on the observation that, if the principal has access to the trajectory of $X$, she can actually compute its quadratic variation $\langle X\rangle$ pathwise. In general, observing the quadratic variation would still prevent the principal from deducing the controls on the volatility, hence the relevance of considering contract of the form (2.11), indexed on $X$ and $\langle X\rangle$ through sensitivity parameters $Z$ and $\Gamma$ to be optimised. The main idea behind our reformulation is to assume further that, since the principal can compute the quadratic variation $\langle X\rangle$, she could actually control it directly. In other words, the idea here is to study an alternative ‘first-best’ formulation of the original principal–agent problem with volatility control, in which the principal enforces a contract $\xi$ as well as a quadratic variation characterised by its density process $\Sigma$. The agent should then optimise his effort in order to maximise his expected utility given such contract $\xi$ as before, but now the admissible set of efforts is restricted to those achieving the quadratic variation imposed by the principal. More precisely, we will consider in the reformulated problem that, given a contract $\xi$ and a quadratic variation density process $\Sigma$, the agent should optimally choose his effort $\nu$ on the drift and the volatility of the output process $X$ so that $\mathrm{d}\langle X\rangle_{t}=\Sigma_{t}\mathrm{d}t$ for all $t\in[0,T]$. Then, the principal’s optimisation problem boils down to choosing the optimal contract $\xi$ and the optimal quadratic variation density process $\Sigma$.
To consider the agent’s problem in this ‘first-best’ reformulation, we shall again fix an admissible contract, namely $\xi\in\Xi$, but also an appropriate process $\Sigma$, representing the density of the quadratic variation with respect to the Lebesgue measure, both would be chosen optimally by the principal, anticipating the agent’s best response. With this in mind, we define the admissible set of processes $\Sigma$ by
@@
For fixed $\Sigma\in{\cal S}$, all probability measures $\mathbb{P}\in{\cal P}^{\circ}(\Sigma)$ are equivalent through Girsanov’s theorem, as mentioned in the previous remark. One can therefore represent the value function of the agent for fixed quadratic variation as a backward stochastic differential equation. Then, since the agent can optimise over admissible quadratic variations, his problem has to be written as the supremum of BSDEs, namely a second-order BSDE. While we will follow the exact same first step, namely to represent the agent’s continuation utility as a BSDE, we will not have to deal with the second step as in our reformulated ‘first-best’ problem, the choice of the quadratic variation is not delegated to the agent, but fixed by the principal.
@@
Recall that by convention, the supremum over an empty set is equal to $\infty$- ∞. In particular, if the principal suggests a quadratic variation which is not implementable by the agent, then the set ${\cal P}^{\circ,\star}(\xi,\Sigma)$ would be empty, leading to a utility of $\infty$- ∞ for the principal, which cannot be optimal. Therefore, we can assume without loss of generality that any potentially optimal $\Sigma\in{\cal S}$ should be such that ${\cal P}^{\circ}(\Sigma)\neq\varnothing$, or alternatively, ${\cal U}^{\circ}(\Sigma)\neq\varnothing$. In other words, for any $\Sigma\in{\cal S}$, there should exist an admissible effort $\nu\in{\cal U}^{\circ}(\Sigma)$, and thus a corresponding probability measure $\mathbb{P}^{\nu}\in{\cal P}^{\circ}(\Sigma)$, such that for all $t\in[0,T]$, $\Sigma_{t}=[\sigma\sigma^{\top}](t,X,\nu_{t}),\;\mathbb{P}^{\nu}$–a.s. This implies in particular that for all $t\in[0,T]$, $\Sigma_{t}$ should take values in $\mathbb{S}_{t}(X)$, defined for all $x\in\Omega$ by
$\mathbb{S}_{t}(x):=\big{\{}S\in\mathbb{S}_{+}^{d}\;:\;S=\big{[}\sigma\sigma^{\top}\big{]}(t,x,u)\;\text{ for some }u\in U\big{\}}.$
Furthermore, recall that the original set ${\cal U}$ of admissible controls is defined such that, for any $\nu\in{\cal U}$, there exists a unique weak solution to (2.3). In particular, the set ${\cal U}^{\circ}(\Sigma)$ of admissible (constrained) controls also satisfies this property. Therefore, under the previous restriction that $\Sigma\in{\cal S}$ should be such that ${\cal U}^{\circ}(\Sigma)\neq\varnothing$, it is clear that weak uniqueness also holds for SDE (3.3). In other words, for fixed $\Sigma\in{\cal S}$, the weak solution $\widebar{\mathbb{P}}$ to SDE (3.3) defined in Remark 3.1 is actually unique, and therefore admits the predictable martingale representation property (see for example Jacod and Shiryaev (2003) [19, Theorem III.4.29]).
@@
Recall that through the original definition of admissible efforts in (2.4), we assume that each $\nu\in{\cal U}$ induces a unique weak solution to (2.3), which implies, as mentioned above, that for fixed $\Sigma\in{\cal S}$, the SDE (3.3) admits a unique weak solution denoted $\widebar{\mathbb{P}}$. This assumption, which is not enforced in [9], is considered here to avoid technical considerations regarding the martingale representation property. Indeed, without weak uniqueness, the solution $\widebar{\mathbb{P}}$ to SDE (3.3) may not admit the predictable martingale representation property. More precisely, the decomposition of any $(\mathbb{F},\widebar{\mathbb{P}})$-local martingale with respect to $X$ would in general involve an extra martingale term, orthogonal to $X$. As this martingale representation property is used to derive the optimal form of contracts, such an additional martingale term would complicate the study. In [9], this additional martingale term can be hidden in the component ‘$K$’ of the solution $(Y,Z,K)$ to the appropriate second-order BSDE, as the set of probability measures considered is naturally saturated (see Possamaï, Tan, and Zhou (2018) [28, Definitions 5.1 and 5.2]). We refer back to Remark 2.2 for a discussion on the reasonableness of this assumption.
@@
We first provide in Section 3.2 a very standard approach, relying on BSDEs theory, to characterise the optimal form of contracts for this ‘first-best’ alternative problem. The result is given in Theorem 3.9. We then prove the main result of this paper, namely Theorem 3.10, which states the equivalence between 1 and 3, i.e. $V_{\rm P}=V_{\rm P}^{\circ}$. Before proceeding, one can already notice that the principal’s value $V_{\rm P}^{\circ}$ in the alternative ‘first-best’ problem summarised above, namely 3, is greater than the value $V_{\rm P}$ in the original problem (1). This result is stated in the following Lemma 3.4, whose proof mostly relies on the observation that, in 3, the principal has an additional control compared to the original problem.
@@
The value of 3 is greater than the value of 1, i.e. $V_{\rm P}^{\circ}\geq V_{\rm P}$.
@@
    Recall that we can assume, without loss of generality, that any potentially optimal contract $\xi\in\Xi$ for 1 should be such that ${\cal P}^{\star}(\xi)\neq\varnothing$, otherwise $V_{\rm P}=-\infty$. One can thus consider $\mathbb{P}^{\star}\in{\cal P}^{\star}(\xi)$, and its corresponding optimal effort process $\nu^{\mathbb{P}^{\star}}\in{\cal U}$, leading to the following representation for the output process $X$:
@@
    Note that the process $\Sigma^{\mathbb{P}^{\star}}$ defined above belongs to the set ${\cal S}$ by definition and we have $\nu^{\mathbb{P}^{\star}}\in{\cal U}^{\circ}(\Sigma^{\star})$, or equivalently, $\mathbb{P}^{\star}\in{\cal P}^{\circ}(\Sigma^{\mathbb{P}^{\star}})$. Additionally, by definition of $\mathbb{P}^{\star}\in{\cal P}^{\star}(\xi)$ in (2.8), we have $V_{\rm A}(\xi)=J_{\rm A}(\xi,\mathbb{P}^{\star})\geq R_{\rm A}$, and therefore

### v2

@@ 3. $(iii)$
Usually, Problem 1 cannot be solved directly: the standard approach is to determine an appropriate form of contracts which simplifies both the agent’s and the principal’s problems, and such that the restriction to this form of contracts is without loss of generality.
Note that the previous setting and assumptions are relatively standard in continuous-time principal–agent problems, and in particular coincide with the framework described in [9]. Indeed, as the goal of this paper is to provide a new approach to solve such principal–agent problems, which in particular requires to show that this approach allows to achieve the same results obtained by Cvitanić, Possamaï, and Touzi (2018) [9], the decision was made to closely follow their original framework. It is nevertheless possible that some assumptions could be relaxed, or at least modified to consider alternative frameworks. For example, in the case of a principal–agent problem with non-continuous output processes (and many agents), one can refer to the recent paper by Hernández Santibáñez (2024) [13] for more general assumptions. It should be noted, however, that the latter focuses on control processes that can be represented through equivalent changes of measures, thus a priori excluding volatility control.
@@ 2.3 The existing approach, via 2BSDEs
In this section, we describe briefly and rather informally the general approach developed by Cvitanić, Possamaï, and Touzi (2018) [9], as this will help in comparing it with our new approach, detailed later in Section 3.
@@
The general approach developed in [9] involves first identifying a class of contracts, offered by the principal, that are revealing, in the sense that the agent’s optimal response to the contract can easily be computed. With this in mind, we first define the appropriate agent’s Hamiltonian,
@@
for any $(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d}$. We then introduce the relevant form of contracts, namely $\xi=Y_{T}^{y_{0},Z,\Gamma}$, where the process $Y^{y_{0},Z,\Gamma}$ is defined as the solution to the following SDE,
@@
for a constant $y_{0}\in\mathbb{R}$ and a pair $(Z,\Gamma)$ of processes satisfying appropriate integrability conditions. More precisely, for any $\mathbb{F}$-predictable $\mathbb{R}^{d}$-valued process $Z$ and any $\mathbb{F}$-optional $\mathbb{R}$-valued process $Y$, we consider the following norms,
@@ 2. $(ii)$
Then, the relevant contracts $\xi\in\Xi$ are those that admit the representation $\xi=Y_{T}^{y_{0},Z,\Gamma}$, where the pair $(Z,\Gamma)\in{\cal V}$ as well as the constant $y_{0}\in\mathbb{R}$ in (2.11) will be optimally chosen by the principal.
@@
Contracts of the form (2.11) are called revealing, as it is straightforward to prove that any agent’s best response to such contracts corresponds to a maximiser of his Hamiltonian. Indeed, relying on Ito’s formula, it is proved in [9, Proposition 3.3] that, given a contract of the form $\xi=Y_{T}^{y_{0},Z,\Gamma}$, $(y_{0},Z,\Gamma)\in\mathbb{R}\times{\cal V}$, any optimal response $\mathbb{P}^{\star}\in{\cal P}^{\star}(\xi)$ is such that
@@
so that the agent’s best response to a contract $\xi=Y_{T}^{y_{0},Z,\Gamma}$ can be written as $\nu^{\mathbb{P}^{\star}}_{t}:=u^{\star}(t,X,Y_{t}^{y_{0},Z,\Gamma},Z_{t},\Gamma_{t})$, $t\in[0,T]$, $\mathbb{P}^{\star}$–a.s. One can also prove that the value of the agent’s optimal control problem associated to such contract $\xi$ of the form (2.11) boils down to $y_{0}$, namely $V_{\rm A}(Y_{T}^{y_{0},Z,\Gamma})=y_{0}$. Consequently, to ensure the agent’s participation in the contractual agreement, it is clear that the principal must choose $y_{0}\geq R_{\rm A}$.
@@
$\mathrm{d}Y_{t}^{y_{0},Z,\Gamma}=$ $\ \Big{(}c_{\rm A}\big{(}t,X,\nu_{t}^{\mathbb{P}^{\star}}\big{)}+Y_{t}^{y_{0},Z,\Gamma}k_{\rm A}\big{(}t,X,\nu_{t}^{\mathbb{P}^{\star}}\big{)}\Big{)}\mathrm{d}t+Z_{t}\cdot\sigma\big{(}t,X,\nu_{t}^{\mathbb{P}^{\star}}\big{)}\mathrm{d}W_{t},\quad t\in[0,T].$ (2.16)
@@
$\mathrm{d}Y_{t}$ $=\Big{(}c_{\rm A}\big{(}t,X,\nu^{\mathbb{P}^{\star}}_{t}\big{)}+Y_{t}\,k_{\rm A}\big{(}t,X,\nu^{\mathbb{P}^{\star}}_{t}\big{)}\Big{)}\mathrm{d}t+Z_{t}\cdot\sigma\big{(}t,X,\nu^{\mathbb{P}^{\star}}_{t}\big{)}\mathrm{d}W_{t},\quad t\in[0,T],$ (2.17b)
coupled through the agent’s optimal effort $\nu^{\mathbb{P}^{\star}}_{t}:=u^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t})$ under $\mathbb{P}^{\star}\in{\cal P}^{\star}(Y_{T})$.
In a Markovian setting, one can solve this stochastic control problem as usual, namely by writing the principal’s Hamiltonian as well as the Hamilton-Jacobi-Bellman (HJB for short) equation satisfied by the value function $\underline{V}_{\rm P}(y_{0})$ (see [9, Theorem 3.9]), and finally optimise it over $y_{0}\geq R_{\rm A}$. More importantly for our purpose in this paper, one can already remark that the value of Problem 2 should be lower than the value in Problem 1, since considering contracts $\xi$ of the form (2.11) is a restriction of the original set $\Xi$ of admissible contracts. This result can be found in [9, Proposition 3.4], and is a direct consequence of [9, Proposition 3.3] mentioned above. For further reference, this result is stated in the lemma below.
@@
The original value defined by Problem 1 is greater than the value in Problem 2, namely $V_{\rm P}\geq\widetilde{V}_{\rm P}$.
@@
To summarise, when restricting the study to contracts $\xi\in\Xi$ of the form (2.11), one can first easily determine the agent’s optimal response to such contract, namely by computing the maximiser of his Hamiltonian. Then, still under this restriction, the principal’s problem boils down to a more standard stochastic problem, namely Problem 2. These two steps are therefore quite straightforward using standard tools from stochastic control theory. However, the last and crucial step requires to prove that the restriction to revealing contracts is without loss of generality. More precisely, one need to argue that the value of the original principal’s problem, namely $V_{\rm P}$ recalled in Problem 1, actually coincides with the value when restricting to revealing contracts, i.e. $\widetilde{V}_{\rm P}$ in Problem 2. While it is clear that $V_{\rm P}\geq\widetilde{V}_{\rm P}$, as mentioned in Lemma 2.5 above, the main mathematical challenge is to show the converse inequality, namely $V_{\rm P}\leq\widetilde{V}_{\rm P}$. This result, stated in [9, Theorem 3.6], relies on the sophisticated theory of second-order BSDEs. We deliberately omit to further develop this step here, as it is precisely the part that our new approach allows us to bypass. It nevertheless seems essential to stress that this is the most difficult stage, which probably prevents a wider application of principal–agent problems with volatility control, but also hampers natural extensions of these problems, as discussed in introduction. Surprisingly, our approach reveals that if Assumption B is not satisfied, the method in [9] may fail to yield optimal contracts. We refer the reader to Remark 3.12 for a more detailed discussion of this point, and to Section 4.3 for an illustrative example.
@@ 3 A new approach
In this section, we first start by formalising a ‘first-best’ alternative problem, in which the principal, in addition to offering a contract, also controls the quadratic variation of the output process. We then provide in Section 3.2 a straightforward resolution approach for this alternative problem, relying on the theory of BSDEs only. We will finally prove in Section 3.3 that this reformulation actually leads to the exact same solution compared to the original problem considered in Section 2.
@@ 3.1 The ‘first-best’ formulation
As already mentioned, the approach developed by Cvitanić, Possamaï, and Touzi (2018) [9] mostly relies on the observation that, if the principal has access to the trajectory of $X$, she can actually compute its quadratic variation $\langle X\rangle$ pathwise. In general, observing the quadratic variation would still prevent the principal from deducing the controls on the volatility, hence the relevance of considering contract of the form (2.11), indexed on $X$ and $\langle X\rangle$ through sensitivity parameters $Z$ and $\Gamma$ to be optimised. The main idea behind our reformulation is to assume further that, since the principal can compute the quadratic variation $\langle X\rangle$, she could actually directly control it. In other words, the idea here is to study an alternative ‘first-best’ formulation of the original principal–agent problem with volatility control, in which the principal enforces a contract $\xi$ as well as a quadratic variation characterised by its density process $\Sigma$. The agent should then optimise his effort in order to maximise his expected utility given such contract $\xi$ as before, but now the admissible set of efforts is restricted to those achieving the quadratic variation imposed by the principal. More precisely, we will consider in the reformulated problem that, given a contract $\xi$ and a quadratic variation density process $\Sigma$, the agent should optimally choose his effort $\nu$ on the drift and the volatility of the output process $X$ so that $\mathrm{d}\langle X\rangle_{t}=\Sigma_{t}\mathrm{d}t$ for all $t\in[0,T]$. Then, the principal’s optimisation problem boils down to choosing the optimal contract $\xi$ and the optimal quadratic variation density process $\Sigma$.
To consider the agent’s problem in this ‘first-best’ reformulation, we shall again fix a contract $\xi$, but also an appropriate process $\Sigma$, representing the density of the quadratic variation with respect to the Lebesgue measure. Both would be chosen optimally by the principal, anticipating the agent’s best response. With this in mind, we define the admissible set of processes $\Sigma$ by
@@
For fixed $\Sigma\in{\cal S}$, all probability measures $\mathbb{P}\in{\cal P}^{\circ}(\Sigma)$ are equivalent through Girsanov’s theorem, as mentioned in the previous remark. One can therefore represent the value function of the agent for fixed quadratic variation as a backward stochastic differential equation. Then, since the agent can optimise over admissible quadratic variations, his problem becomes a supremum of BSDEs, i.e. a second-order BSDE. While we will follow the exact same first step, namely to represent the agent’s continuation utility as a BSDE, we will not have to deal with the second step as in our ‘first-best’ problem, the quadratic variation is chosen by the principal, not delegated to the agent.
@@
Recall that by convention, the supremum over an empty set is equal to $\infty$- ∞. In particular, if the principal suggests a quadratic variation which is not implementable by the agent, then the set ${\cal P}^{\circ,\star}(\xi,\Sigma)$ would be empty, leading to a utility of $\infty$- ∞ for the principal, which cannot be optimal. Therefore, we can assume without loss of generality that for any potentially optimal $\Sigma\in{\cal S}$, there exists an admissible effort $\nu\in{\cal U}^{\circ}(\Sigma)$, and thus a corresponding probability measure $\mathbb{P}^{\nu}\in{\cal P}^{\circ}(\Sigma)$, such that for all $t\in[0,T]$, $\Sigma_{t}=[\sigma\sigma^{\top}](t,X,\nu_{t}),\;\mathbb{P}^{\nu}$–a.s. This implies in particular that for all $t\in[0,T]$, $\Sigma_{t}$ should take values in $\mathbb{S}_{t}(X)$, defined for all $x\in\Omega$ by
$\mathbb{S}_{t}(x):=\big{\{}S\in\mathbb{S}_{+}^{d}\;:\;S=\big{[}\sigma\sigma^{\top}\big{]}(t,x,u)\;\text{ for some }u\in U\big{\}}.$ (3.9)
Furthermore, recall that the original set ${\cal U}$ of admissible controls is defined in Assumption A such that, for any control $\nu\in{\cal U}$, there exists a unique weak solution to the controlled SDE (2.3). In particular, the set ${\cal U}^{\circ}(\Sigma)$ of admissible (constrained) controls also satisfies this property. Therefore, under the previous restriction that $\Sigma\in{\cal S}$ should be such that ${\cal U}^{\circ}(\Sigma)\neq\varnothing$, it is clear that weak uniqueness also holds for SDE (3.3). In other words, for fixed $\Sigma\in{\cal S}$, the weak solution $\widebar{\mathbb{P}}$ to SDE (3.3) defined in Remark 3.1 is actually unique, and therefore admits the predictable martingale representation property (see for example Jacod and Shiryaev (2003) [19, Theorem III.4.29]).
@@
This assumption that weak uniqueness holds for SDE (2.3), and thus for SDE (3.3), which is not enforced in [9], is considered here to avoid technical considerations regarding the martingale representation property. Indeed, without weak uniqueness, the solution $\widebar{\mathbb{P}}$ to SDE (3.3) may not admit the predictable martingale representation property. More precisely, the decomposition of any $(\mathbb{F},\widebar{\mathbb{P}})$-local martingale with respect to $X$ would in general involve an extra martingale term, orthogonal to $X$. As this martingale representation property is used to derive the optimal form of contracts, such an additional martingale term would complicate the study. In [9], this additional martingale term can be hidden in the component ‘$K$’ of the solution $(Y,Z,K)$ to the appropriate second-order BSDE, as the set of probability measures considered is naturally saturated (see Possamaï, Tan, and Zhou (2018) [28, Definitions 5.1 and 5.2]). We refer back to Remark 2.2 for a discussion on the reasonableness of this assumption.
@@
We first provide in Section 3.2 a very standard approach, relying on BSDEs theory, to characterise the optimal form of contracts for this ‘first-best’ alternative problem. The result is given in Theorem 3.9. We then prove the main result of this paper, namely Theorem 3.10, which states the equivalence between Problem 1 and Problem 3, i.e. $V_{\rm P}=V_{\rm P}^{\circ}$. Before proceeding, one can already notice that the principal’s value $V_{\rm P}^{\circ}$ in the alternative ‘first-best’ problem summarised above, namely Problem 3, is greater than the value $V_{\rm P}$ in the original problem (Problem 1). This result is stated in the following Lemma 3.4, whose proof mostly relies on the observation that, in Problem 3, the principal has an additional control compared to the original problem.
@@
The value of Problem 3 is greater than the value of Problem 1, i.e. $V_{\rm P}^{\circ}\geq V_{\rm P}$.
@@
    Recall that we can assume, without loss of generality, that any potentially optimal contract $\xi\in\Xi$ for Problem 1 should be such that ${\cal P}^{\star}(\xi)\neq\varnothing$, otherwise $V_{\rm P}=-\infty$. One can thus consider $\mathbb{P}^{\star}\in{\cal P}^{\star}(\xi)$, and its corresponding optimal effort process $\nu^{\mathbb{P}^{\star}}\in{\cal U}$, leading to the following representation for the output process $X$:
@@
    Note that the process $\Sigma^{\mathbb{P}^{\star}}$ defined above belongs to the set ${\cal S}$ by definition and we have $\nu^{\mathbb{P}^{\star}}\in{\cal U}^{\circ}(\Sigma^{\mathbb{P}^{\star}})$, or equivalently, $\mathbb{P}^{\star}\in{\cal P}^{\circ}(\Sigma^{\mathbb{P}^{\star}})$. Additionally, by definition of $\mathbb{P}^{\star}\in{\cal P}^{\star}(\xi)$ in (2.8), we have $V_{\rm A}(\xi)=J_{\rm A}(\xi,\mathbb{P}^{\star})\geq R_{\rm A}$, and therefore